OpenStudy (anonymous):

Solve the system of equations: y = 0.5x + 2 -y = -2x + 4

hero (hero):

you can essentially do the same thing with this one

hero (hero):

But, you might want to just do it this way: y = 0.5x + 2 y = 2x - 4 y = y 0.5x + 2 = 2x - 4 1.5x = 6 x = 6/1.5 x = 4 y = 4

OpenStudy (anonymous):

Thank You